# Extending Douglas–Rachford Splitting for Convex Optimization

### v1

@@ 1 Introduction
The uniqueness of the Douglas–Rachford method under these conditions for problems of the form (1) is established in [ryu2020uniqueness] through two results: one parameterizing the class of frugal, no-lifting resolvent methods with the fixed-point encoding property, and one characterizing its subset of methods with unconditional convergence. It was shown that each frugal, no-lifting resolvent method for solving (1)—up to equivalence—satisfies the fixed-point encoding property if and only if it has the form
@@
$S^{(2)}$ $\coloneqq\left\{(\alpha,\beta,\theta)\in\mathbb{R}_{++}^{3}\mid\alpha/\beta\in[1,\infty),\ \theta\in(0,2)\right\},$
@@
While we primarily characterize the entire class of frugal, minimal lifting, unconditionally convergent resolvent splittings with fixed-point encoding in the optimization setting, we emphasize that all our convergence guarantees for methods with $\alpha/\beta\neq 1$ appear to be new. Because the alternating direction method of multipliers (ADMM) [GlowinskiMarroco1975, GabayMercier1976, BoydParikhChuPeleatoEckstein2011] arises as a special case of Douglas–Rachford splitting [Gabay1983, EcksteinBertsekas1992], our results yield a family of completely new convergent ADMM-type methods for convex optimization. Likewise, since the Chambolle–Pock algorithm [chambolle2011first] is also a Douglas–Rachford instance [o2020equivalence], we obtain an expanded class of convergent Chambolle–Pock–type methods, some of which are new. In particular, we present the first doubly relaxed Chambolle–Pock-type method (see Algorithm˜4 for the specific relaxations) with freedom in both relaxation parameters. Previous works, such as [he2012convergence, banert2025chambolle], have one of them fixed.
The rest of the paper is organized as follows. Section˜2 introduces notation and definitions and presents a few preliminary results. Section˜3 presents the problem setting and our main result; the complete characterization of all frugal, no-lifting resolvent splittings in the convex optimization setting that are both unconditionally convergent and has the fixed-point encoding property. Section˜4 is devoted to proving the main result. Section˜5 shows ergodic $\mathcal{O}(1/K)$-convergence for two primal-dual gap functions for the characterized methods. Section˜6 applies our method to specific problem instances to derive new convergent ADMM-, Chambolle–Pock-, and parallel splitting-type algorithms. Section˜7 presents an operator composition perspective of the algorithm in (3) and Section˜8 concludes the paper.
@@ 2.1 Notation
Our notation closely follows [rockafellar1970convex, BauschkeCombettes2017]. We let $\mathbb{N}=\{0,1,\dots\}$ denote the natural numbers, $\mathbb{R}$ denote the real numbers, $\mathbb{R}_{+}$ the nonnegative numbers, and $\mathbb{R}_{++}$ the positive numbers. Moreover $\mathbb{R}^{n}$ denotes the $n$-dimensional real vectors, $\mathbb{R}^{n\times m}$ denotes the real $n\times m$ matrices, and $\mathbb{S}^{n}$ the real symmetric $n\times n$ matrices.
@@
for each $x\in\mathcal{H}$. The subdifferential $\partial f$ is maximally monotone [BauschkeCombettes2017, Theorem 20.25], and its resolvent coincides with the proximal operator of $f$, i.e., ${\rm{prox}}_{f}\coloneqq J_{\partial f}$; see [BauschkeCombettes2017, Chapter 24] for details. For a function $f:\mathcal{H}\rightarrow\mathbb{R}\cup\{+\infty\}$ we let $\operatorname*{Argmin}_{x\in\mathcal{H}}(f(x))=\{x\in\mathcal{H}\mid f(x)=\inf_{z\in\mathcal{H}}f(z)\}$
Let $C$ be a subset of $\mathcal{H}$. We let $\iota_{C}:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ denote the indicator function of $C$, which equals 0 for $x\in C$ and $\infty$ for $x\not\in C$.
@@
$\partial\mathcal{F}\coloneqq\{A:\mathcal{H}\rightrightarrows\mathcal{H}\mid A=\partial f\text{ with }f\in\mathcal{F}\}.$
@@ 3 Problem Setting and Main Results
In this section, we present our main result: a complete characterization of the frugal, no-lifting resolvent-splitting methods that are unconditionally convergent and satisfy the fixed-point encoding property for problems of the form (1) with $A=\partial f$ and $B=\partial g$, where $f,g:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ are proper, closed, and convex. The starting point for our analysis is Theorem 1 of [ryu2020uniqueness], which gives an explicit parameterization of all frugal, no-lifting resolvent-splitting methods satisfying the fixed-point encoding property, under the assumption that the two operators are maximally monotone. Since the proof of that result relies only on subdifferential operators of proper, closed, convex functions, the same parameterization applies in our setting. This parameterization is given explicitly by Algorithm˜1, and hence the task of obtaining our characterization reduces to determining which choices of $(\alpha,\beta,\theta)$ in Algorithm˜1 lead to unconditional convergence in the convex optimization setting.
@@
Let $f,g:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ be proper, closed, and convex. Up to equivalence, a method is a frugal, no-lifting resolvent-splitting satisfying the fixed-point encoding property for solving ˜1 with $A=\partial f$ and $B=\partial g$ if and only if it is of the form Algorithm˜1.
@@
$S^{(2)}$ $\coloneqq\left\{(\alpha,\beta,\theta)\in\mathbb{R}_{++}^{3}\mid\alpha/\beta\in[1,\infty),\ \theta\in(0,2)\right\}.$
@@
In contrast to the maximal monotone setting, no additional assumptions on the dimension of $\mathcal{H}$ are needed when restricting to $\partial\mathcal{F}\times\partial\mathcal{F}$ or $\partial\mathcal{F}\times\mathcal{A}$. In particular, Theorem˜3.3 also covers the one-dimensional case. In light of the fact that every maximally monotone operator on $\mathbb{R}$ is the subdifferential of a proper, closed, and convex function (see [BauschkeCombettes2017, Corollary 22.23]) and since any one-dimensional $\mathcal{H}$ is isometrically isomorphic to $\mathbb{R}$ which preserves maximal monotonicity and subdifferentials, Theorem˜3.3 allows us to revisit [ryu2020uniqueness, Theorem 2] and close the gap left open in the case $\dim\mathcal{H}=1$.
@@ 4 Convergence Analysis
In Section˜4.2, we prove necessity by constructing explicit counterexamples that do not converge outside these regions, thereby showing that
@@ 4.1 Lyapunov Analysis
where $f\in\mathcal{F}$ and $B\in\mathcal{A}$. This includes inclusion problems over $\partial\mathcal{F}\times\partial\mathcal{F}$ as a special case.
We will consider sequences $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$, generated by Algorithm˜1 applied to $(\partial f,B,\alpha,\beta,\theta)\in\partial\mathcal{F}\times\mathcal{A}\times\mathbb{R}_{++}^{2}\times(\mathbb{R}\setminus\{0\})$. We first introduce some auxiliary variables in the following definition to simplify notation.
Definition 4.1 ($u^{\star},u^{k}_{i}$):
Let $(f,B)\in\mathcal{F}\times\mathcal{A}$ such that $\operatorname*{zer}(\partial f+B)\neq\emptyset$ and $(\alpha,\beta,\theta)\in\mathbb{R}^{2}_{++}\times(\mathbb{R}\setminus\{0\})$. Let $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ be generated by Algorithm˜1 applied to $(\partial f,B,\alpha,\beta,\theta)$. Let $(x^{\star},x^{\star},z^{\star})\in\operatorname{Fix}(\partial f,B,\alpha,\beta)$. Then, define
$u^{\star}\coloneqq\frac{1}{\alpha}(z^{\star}-x^{\star})\in\partial f(x^{\star})$
and define the sequences $(u_{1}^{k})_{k\in\mathbb{N}}$ and $(u_{2}^{k})_{k\in\mathbb{N}}$ by
$u_{1}^{k}$ $\coloneqq\frac{1}{\alpha}(z^{k}-x^{k}_{1})\in\partial f(x^{k}_{1})$
$u^{k}_{2}$ $\coloneqq\frac{1}{\beta}\left(\left(1+\frac{\beta}{\alpha}\right)x^{k}_{1}-\frac{\beta}{\alpha}z^{k}-x^{k}_{2}\right)\in B(x^{k}_{2})$
for every $k\in\mathbb{N}$.
We now introduce the Lyapunov function $V_{k}$ and the residual function $R_{k}$. Since the analysis separates into the two parameter regions $S^{(1)}$ and $S^{(2)}$, we define two versions $V_{k}^{(i)}$ and $R_{k}^{(i)}$ for $i\in\{1,2\}$, each tailored to its respective region. When $(\alpha,\beta,\theta)\in S^{(i)}$, the pair $\left(V_{k}^{(i)},R_{k}^{(i)}\right)$ is the one used in the convergence analysis.
[…]
Let $(f,B)\in\mathcal{F}\times\mathcal{A}$ such that $\operatorname*{zer}(\partial f+B)\neq\emptyset$ and $(\alpha,\beta,\theta)\in\mathbb{R}^{2}_{++}\times(\mathbb{R}\setminus\{0\})$ such that $\theta\neq 4\alpha/\beta$. Let $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ be generated by Algorithm˜1 applied to $(\partial f,B,\alpha,\beta,\theta)$. Let $(x^{\star},x^{\star},z^{\star})\in\operatorname{Fix}(\partial f,B,\alpha,\beta)$ and $\tau\coloneqq\alpha/\beta\in\mathbb{R}_{++}$. Then, define
@@
$V^{(1)}_{k}$ $\coloneqq Q^{(1)}_{k}+2\alpha(1-\tau)D_{f}(x^{k}_{1},x^{\star},u^{\star}),$
$R^{(1)}_{k}$ $\coloneqq\theta(2\tau-\theta)\left\lVert x_{1}^{k}-x_{2}^{k}\right\rVert^{2}+(1-\tau)\left\lVert x_{1}^{k+1}-x_{1}^{k}\right\rVert^{2}+2\alpha(1-\tau)D_{f}(x^{k}_{1},x^{k+1}_{1},u^{k+1}_{1})$
and
@@
$\quad\;+\frac{\tau(\tau-1)(4-\theta)}{4\tau-\theta}\left\lVert x_{1}^{k}-x^{\star}\right\rVert^{2},$
$V^{(2)}_{k}$ $\coloneqq Q^{(2)}_{k}+2\alpha(\tau-1)D_{f}(x^{\star},x^{k}_{1},u^{k}_{1}),$
$R^{(2)}_{k}$ $\coloneqq(\tau-1)\left\lVert x_{1}^{k+1}-x_{1}^{k}+\theta(x_{1}^{k}-x_{2}^{k})\right\rVert^{2}+\tau\theta(2-\theta)\left\lVert x_{1}^{k}-x_{2}^{k}\right\rVert^{2}+2\alpha(\tau-1)D_{f}(x^{k+1}_{1},x^{k}_{1},u^{k}_{1}),$
@@
These Lyapunov and residual functions are in fact the same on the region we consider, i.e., $V_{k}^{(1)}=V_{k}^{(2)}$ and $R_{k}^{(1)}=R_{k}^{(2)}$ for all $(\alpha,\beta,\theta)\in S^{(1)}\cup S^{(2)}$. We nevertheless use these two representations since it, for instance, is not immediately clear why $V_{k}^{(1)}$ and $R_{k}^{(1)}$ are nonnegative for $(\alpha,\beta,\theta)\in S^{(2)}$ and conversely for $V_{k}^{(2)}$ and $R_{k}^{(2)}$ on $(\alpha,\beta,\theta)\in S^{(1)}$. Working with both representations makes the analysis more transparent. The identities $V_{k}^{(1)}=V_{k}^{(2)}$ and $R_{k}^{(1)}=R_{k}^{(2)}$ are verified symbolically in [nilsson2025verification], but the analysis to come does not rely on these identities.
Remark 4.2:
Note that if $(\alpha,\beta,\theta)\in S^{(1)}\cup S^{(2)}$, then $4\alpha/\beta>\theta$. This implies, in particular, that the definition of $Q_{k}^{(2)}$ never involves division by 0.
Remark 4.3:
Note that, since $\operatorname*{ran}J_{\alpha\partial f}=\operatorname*{dom}\partial f\subset\operatorname*{dom}f$, by [BauschkeCombettes2017, Proposition 23.2(i), Proposition 16.4(i)], the Bregman-type distances $D_{f}$ in $V^{(i)}_{k}$ and $R^{(i)}_{k}$ are well-defined and nonnegative according to Definitions˜2.1, 2.3 and 4.1 for all $i\in\{1,2\}$ and all $k\in\mathbb{N}$.
Remark 4.4:
Note that there for each $i\in\{1,2\}$ exists a unique matrix $Q_{V}^{(i)}\in\mathbb{S}^{3}$ such that
$Q^{(i)}_{k}=\mathcal{Q}(Q_{V}^{(i)},(x^{k}_{1}-x^{\star},x^{k}_{2}-x^{\star},z^{k}-z^{\star}))$
holds for all $k\in\mathbb{N}$ and all $(x^{\star},x^{\star},z^{\star})\in\operatorname{Fix}(\partial f,B,\alpha,\beta)$. Viewing $Q^{(i)}_{k}$ as a quadratic form in this way will simplify the proof of Proposition˜4.3, see Appendix˜A for the explicit expression for $Q_{V}^{(i)}$, lifted to $\mathbb{R}^{5}\times\mathbb{R}^{5}$.
@@
With $u_{i}^{k},u^{\star}\in\mathcal{H}$ as in Definition˜4.1, let
@@
The quantity $I_{k}$ in Definition˜4.3 satisfies $I_{k}\geq 0$ for all $k\in\mathbb{N}$.
@@
Let $k\in\mathbb{N}$. By Definition˜4.1 we get that each of the four inner products in $I_{k}$ is a monotonicity term; the first and third arise from the monotonicity of $\partial f$ evaluated at $(x_{1}^{k},x^{\star})$ and $(x_{1}^{k+1},x^{\star})$, while the second and fourth arise from the monotonicity of $B$ evaluated at $(x_{2}^{k},x^{\star})$ and $(x_{2}^{k+1},x^{\star})$. Therefore, $I_{k}\geq 0$. ∎
The following Lyapunov equality is the key component for establishing unconditional convergence of Algorithm˜1.
@@
With $V_{k}^{(i)}$ and $R_{k}^{(i)}$ as in Definition˜4.2 and $I_{k}$ as in Definition˜4.3, if $(\alpha,\beta,\theta)\in S^{(i)}$, then $V_{k}^{(i)}\geq 0$, $R_{k}^{(i)}\geq 0$, and
$V_{k+1}^{(i)}=V_{k}^{(i)}-R_{k}^{(i)}-\theta\alpha I_{k},$ (7)
holds for all $k\in\mathbb{N}$ and $i\in\{1,2\}$.
@@
The fact that $V_{k}^{(i)}$ and $R_{k}^{(i)}$ for both $i\in\{1,2\}$ are nonnegative follows immediately from their respective definitions, since all coefficients in front of the norms and the Bregman-type distances are nonnegative when $(\alpha,\beta,\theta)\in S^{(i)}$ and that the Bregman-type distances are nonnegative by Lemma˜2.3. The equality follows after expanding every norm expression of $V_{k}^{(i)},V_{k+1}^{(i)}$, and $R_{k}^{(i)}$, using the update step $z^{k+1}=z^{k}+\theta(x_{2}^{k}-x_{1}^{k})$, using Lemma˜2.1, and verifying that
$V_{k}^{(i)}-V_{k+1}^{(i)}-R_{k}^{(i)}=\theta\alpha I_{k}$ (LE)
holds for all $k\in\mathbb{N}$ and $i\in\{1,2\}$. The Lyapunov equality in (LE) is verified for both $i\in\{1,2\}$ in Appendix˜A and symbolically in [nilsson2025verification]. ∎
@@
With $(x^{k}_{i})_{k\in\mathbb{N}}$ and $(z^{k})_{k\in\mathbb{N}}$ generated by Algorithm˜1, $(u^{k}_{i})_{k\in\mathbb{N}}$ as in Definition˜4.1, $V_{k}^{(i)},Q_{k}^{(i)},R_{k}^{(i)}$ as in Definition˜4.2, and $I_{k}$ as in Definition˜4.3, we have that if $(\alpha,\beta,\theta)\in S^{(i)}$, then
@@
    $\left(R_{k}^{(i)}\right)_{k\in\mathbb{N}}$ and $\left(I_{k}\right)_{k\in\mathbb{N}}$ are summable,
@@
    $\left(Q_{k}^{(i)}\right)_{k\in\mathbb{N}}$ converges,
holds for all $i\in\{1,2\}$.

### v2

@@ 1 Introduction
The uniqueness of the Douglas–Rachford method under these conditions for problems of the form (1) is established in [ryu2020uniqueness] through two results: one parameterizing the class of frugal, no-lifting resolvent methods with the fixed-point encoding property, and one characterizing its subset of methods with unconditional convergence. It was shown that each frugal, no-lifting resolvent method for solving (1)—up to equivalence—satisfies the fixed-point encoding property if and only if it is of the form
@@
$S^{(2)}$ $\coloneqq\left\{(\alpha,\beta,\theta)\in\mathbb{R}_{++}^{3}\mid\alpha/\beta\in[1,+\infty),\ \theta\in(0,2)\right\},$
@@
While we primarily characterize the entire class of frugal, minimal lifting, unconditionally convergent resolvent splittings with fixed-point encoding in the optimization setting, we emphasize that all our convergence guarantees for methods with $\alpha/\beta\neq 1$ appear to be new. Because the alternating direction method of multipliers (ADMM) [GlowinskiMarroco1975, GabayMercier1976, BoydParikhChuPeleatoEckstein2011] arises as a special case of Douglas–Rachford splitting [Gabay1983, EcksteinBertsekas1992], our results yield a family of completely new convergent ADMM-type methods for convex optimization. Likewise, since the Chambolle–Pock algorithm [chambolle2011first] is also a Douglas–Rachford instance [o2020equivalence], we obtain an expanded class of convergent Chambolle–Pock–type methods, some of which are new. In particular, we present the, to the authors’ knowledge, first doubly relaxed Chambolle–Pock-type method (see Algorithm˜4 for the specific relaxations) with freedom in both relaxation parameters. Previous works, such as [he2012convergence, banert2025chambolle], have one of them fixed. We also present extensions of the parallel splitting method in [BauschkeCombettes2017, Proposition 28.7] and a generalized alternating projections method [GAP_Agmon, GAP_Motzkin, GAP_Bregman].
The rest of the paper is organized as follows. Section˜2 introduces notation and definitions and presents a few preliminary results. Section˜3 presents the problem setting and our main result; the complete characterization of all frugal, no-lifting resolvent splittings in the convex optimization setting that are both unconditionally convergent and have the fixed-point encoding property. Section˜4 is devoted to proving the main result. Section˜5 shows ergodic $\mathcal{O}(1/K)$-convergence for two primal-dual gap functions for the characterized methods. Section˜6 applies our method to specific problem instances to derive new convergent ADMM-, Chambolle–Pock-, parallel splitting-, and alternating projection-type algorithms, and Section˜7 concludes the paper.
@@ 2.1 Notation
Our notation closely follows [rockafellar1970convex, BauschkeCombettes2017]. We let $\mathbb{N}=\{0,1,\dots\}$ denote the natural numbers, $\mathbb{R}$ denote the real numbers, $\mathbb{R}_{+}$ the nonnegative numbers, and $\mathbb{R}_{++}$ the positive numbers. Moreover $\mathbb{R}^{n}$ denotes the $n$-dimensional real vectors, $\mathbb{R}^{n\times m}$ denotes the real $n\times m$ matrices, and $\mathbb{S}^{n}$ denotes the real symmetric $n\times n$ matrices.
@@
for each $x\in\mathcal{H}$. The subdifferential $\partial f$ is maximally monotone [BauschkeCombettes2017, Theorem 20.25], and its resolvent coincides with the proximal operator of $f$, i.e., ${\rm{prox}}_{f}=J_{\partial f}$; see [BauschkeCombettes2017, Chapter 24] for details.
Let $C$ be a subset of $\mathcal{H}$. We let $\iota_{C}:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ denote the indicator function of $C$, which equals 0 for $x\in C$ and $\infty$ for $x\not\in C$, and let $\Pi_{C}:={\rm{prox}}_{\iota_{C}}$ denote the orthogonal projection onto $C$.
@@
$\partial\mathcal{F}\coloneqq\{\partial f:\mathcal{H}\rightrightarrows\mathcal{H}\mid f\in\mathcal{F}\}.$
@@ 3 Problem Setting and Main Results
In this section, we present our main result: a complete characterization of the frugal, no-lifting resolvent-splitting methods that are unconditionally convergent and satisfy the fixed-point encoding property for problems of the form (1) with $A=\partial f$ and $B=\partial g$, where $f,g:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ are proper, closed, and convex functions. The starting point for our analysis is Theorem 1 of [ryu2020uniqueness], which gives an explicit parameterization of all frugal, no-lifting resolvent-splitting methods satisfying the fixed-point encoding property, under the assumption that the two operators are maximally monotone. Since the proof of that result relies only on subdifferential operators of proper, closed, convex functions, the same parameterization applies in our setting. This parameterization is given explicitly by Algorithm˜1, and hence the task of obtaining our characterization reduces to determining which choices of $(\alpha,\beta,\theta)$ in Algorithm˜1 lead to unconditional convergence in the convex optimization setting.
@@
Let $f,g:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ be proper, closed, and convex functions. Up to equivalence, a method is a frugal, no-lifting resolvent-splitting satisfying the fixed-point encoding property for solving ˜1 with $A=\partial f$ and $B=\partial g$ if and only if it is of the form given in Algorithm˜1.
@@
$S^{(2)}$ $\coloneqq\left\{(\alpha,\beta,\theta)\in\mathbb{R}_{++}^{3}\mid\alpha/\beta\in[1,+\infty),\ \theta\in(0,2)\right\}.$
@@
In contrast to the maximal monotone setting, no additional assumptions on the dimension of $\mathcal{H}$ are needed when restricting to $\partial\mathcal{F}\times\partial\mathcal{F}$ or $\partial\mathcal{F}\times\mathcal{A}$. In particular, Theorem˜3.3 also covers the one-dimensional case. In light of the fact that every maximally monotone operator on $\mathbb{R}$ is the subdifferential of a proper, closed, and convex function (see [BauschkeCombettes2017, Corollary 22.23]) and since any one-dimensional $\mathcal{H}$ is isometrically isomorphic to $\mathbb{R}$, with a mapping that preserves maximal monotonicity and subdifferentials, Theorem˜3.3 allows us to revisit [ryu2020uniqueness, Theorem 2] and close the gap left open in the case $\dim\mathcal{H}=1$.
@@
Before we proceed to prove Theorem˜3.3, we note that the iteration in Algorithm˜1 has been analyzed before. In particular, in [DaoPhan2019AdaptiveDRS] it was written as
$z^{k+1}=T_{A,B}^{\alpha,\beta,\theta}(z^{k})\qquad\quad\text{where}\quad\qquad T_{A,B}^{\alpha,\beta,\theta}\coloneqq\left(1-\frac{\theta\beta}{\alpha+\beta}\right)\operatorname*{Id}+\frac{\theta\beta}{\alpha+\beta}R_{B}^{\beta,\alpha}R_{A}^{\alpha,\beta},$ (7)
and $R_{A}^{\alpha,\beta}:\mathcal{H}\to\mathcal{H}$ and $R_{B}^{\beta,\alpha}:\mathcal{H}\to\mathcal{H}$ are relaxed resolvents defined as
$R_{A}^{\alpha,\beta}\coloneqq\left(1+\frac{\beta}{\alpha}\right)J_{\alpha A}-\frac{\beta}{\alpha}\operatorname*{Id}\qquad\quad{\text{and}}\qquad\quad R_{B}^{\beta,\alpha}\coloneqq\left(1+\frac{\alpha}{\beta}\right)J_{\beta B}-\frac{\alpha}{\beta}\operatorname*{Id}.$ (8)
This formulation has been analyzed, e.g., in [DaoPhan2019AdaptiveDRS, GiselssonMoursi2021Compositions, BartzDaoPhan2022ConicalAveragedness, Dao_2025] in different settings than ours. Under the parameter restrictions in Theorem˜3.3, we have $\frac{\theta\beta}{\alpha+\beta}\in(0,1)$, so (7) is a $\frac{\theta\beta}{\alpha+\beta}$-averaged iteration of the composition $R_{B}^{\beta,\alpha}R_{A}^{\alpha,\beta}$. It turns out that $T_{A,B}^{\alpha,\beta,\theta}$ is an averaged operator [BauschkeCombettes2017, Definition 4.33] for all $(A,B)\in\mathcal{A}\times\mathcal{A}$ if and only if $\alpha=\beta$—in which case the two relaxed resolvents are nonexpansive—and it can fail to be averaged even for the subclass $\partial\mathcal{F}\times\partial\mathcal{F}$ unless $\alpha=\beta$. Consequently, in contrast to the standard Douglas–Rachford case, the proof of Theorem˜3.3 cannot rely on averaged-operator theory and instead makes use of a Lyapunov analysis, as developed in Section˜4.
@@ 4 Convergence Analysis
In Section˜4.2, we prove necessity by constructing explicit counterexamples for which the algorithm does not converge outside these regions, thereby showing that
@@ 4.1 Lyapunov Analysis
where $f\in\mathcal{F}$ and $B\in\mathcal{A}$. This includes inclusion problems over $\partial\mathcal{F}\times\partial\mathcal{F}$ as a special case. Throughout Section˜4.1 we will assume the following.
Assumption 4.1:
Suppose that
  1. (i)
    $(f,B)\in\mathcal{F}\times\mathcal{A}$ such that $\operatorname*{zer}(\partial f+B)\neq\emptyset$,
  2. (ii)
    $(\alpha,\beta,\theta)\in\mathbb{R}_{++}^{2}\times(\mathbb{R}\setminus\{0\})$.
Let us also define the sequences and points needed for the upcoming results.
Definition 4.1:
Under ˜4.1,
  1. (i)
    let $(x^{\star},x^{\star},z^{\star})\in\operatorname{Fix}(\partial f,B,\alpha,\beta)$, where $\operatorname{Fix}(\partial f,B,\alpha,\beta)$ is nonempty by Proposition˜3.1 and ˜4.1LABEL:*ass::Lyapunov_assumption_i, and let $u^{\star}\coloneqq\frac{1}{\alpha}(z^{\star}-x^{\star})$,
  2. (ii)
    let the sequence $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ be generated by Algorithm˜1 applied to $(\partial f,B,z^{0},\alpha,\beta,\theta)$ for some initial point $z^{0}\in\mathcal{H}$,
  3. (iii)
    and let $(u_{1}^{k})_{k\in\mathbb{N}}$ and $(u_{2}^{k})_{k\in\mathbb{N}}$ be sequences such that
    $u_{1}^{k}$ $=\frac{1}{\alpha}(z^{k}-x^{k}_{1}),$
    $u^{k}_{2}$ $=\frac{1}{\beta}\left(\left(1+\frac{\beta}{\alpha}\right)x^{k}_{1}-\frac{\beta}{\alpha}z^{k}-x^{k}_{2}\right),$
    for all $k\in\mathbb{N}$.
Remark 4.1:
For the sequences of Definition˜4.1, we have that $u_{1}^{k}\in\partial f(x_{1}^{k})$ and $u_{2}^{k}\in B(x_{2}^{k})$ for all $k\in\mathbb{N}$, due to the algorithm update of Algorithm˜1. We additionally have that $u^{\star}\in\partial f(x^{\star})$ and $u^{\star}\in B(x^{\star})$.
Next, we introduce the Lyapunov function $V_{k}$ and the residual function $R_{k}$, both of which contain Bregman-type distance terms $D_{f}$. Since $\operatorname*{ran}J_{\alpha\partial f}=\operatorname*{dom}\partial f\subset\operatorname*{dom}f$, by [BauschkeCombettes2017, Proposition 23.2(i), Proposition 16.4(i)], these Bregman-type distances are well-defined and nonnegative for all $k\in\mathbb{N}$ according to Definitions˜2.1, 2.3 and 4.1.
Definition 4.2 ($Q^{(1)}_{k},V_{k},R_{k}$):
For the sequences of Definition˜4.1 and for all $k\in\mathbb{N}$, let $\tau\coloneqq\alpha/\beta$ and define
@@
$V_{k}$ $\coloneqq Q^{(1)}_{k}+2\alpha(1-\tau)D_{f}(x^{k}_{1},x^{\star},u^{\star}),$
$R_{k}$ $\coloneqq\theta(2\tau-\theta)\left\lVert x_{1}^{k}-x_{2}^{k}\right\rVert^{2}+(1-\tau)\left\lVert x_{1}^{k+1}-x_{1}^{k}\right\rVert^{2}+2\alpha(1-\tau)D_{f}(x^{k}_{1},x^{k+1}_{1},u^{k+1}_{1}).$
Nonnegativity of $V_{k}$ and $R_{k}$ for all $(\alpha,\beta,\theta)\in S^{(1)}$ is immediate. The following lemma will be used in Proposition˜4.1 to show that $V_{k}$ and $R_{k}$ are nonnegative for $(\alpha,\beta,\theta)\in S^{(2)}$.
Lemma 4.1:
For the sequences of Definition˜4.1 and for all $k\in\mathbb{N}$, let $\tau\coloneqq\alpha/\beta$ and let
@@
$\quad\;+\frac{\tau(\tau-1)(4-\theta)}{4\tau-\theta}\left\lVert x_{1}^{k}-x^{\star}\right\rVert^{2}.$
If $(\alpha,\beta,\theta)\in S^{(1)}\cup S^{(2)}$, then $V_{k}$ and $R_{k}$ satisfy
$V_{k}$ $=Q^{(2)}_{k}+2\alpha(\tau-1)D_{f}(x^{\star},x^{k}_{1},u^{k}_{1}),$ (VR)
$R_{k}$ $=(\tau-1)\left\lVert x_{1}^{k+1}-x_{1}^{k}+\theta(x_{1}^{k}-x_{2}^{k})\right\rVert^{2}+\tau\theta(2-\theta)\left\lVert x_{1}^{k}-x_{2}^{k}\right\rVert^{2}+2\alpha(\tau-1)D_{f}(x^{k+1}_{1},x^{k}_{1},u^{k}_{1})$
@@
Proof.
Note first that if $(\alpha,\beta,\theta)\in S^{(1)}\cup S^{(2)}$, then $4\alpha/\beta>\theta$. This implies, in particular, that the definition of $Q_{k}^{(2)}$ never involves division by 0. The equalities (VR) are shown in Section˜A.1 and symbolically in [^1] https://github.com/MaxNilsson-phd/Symbolic_verification_of_Extending_Douglas--Rachford_Splitting_for_Convex_Optimization. ∎
@@
Note that for each $i\in\{1,2\}$ there exists a unique matrix $\bar{Q}_{V}^{(i)}\in\mathbb{S}^{3}$ such that the equality
$Q^{(i)}_{k}=\mathcal{Q}(\bar{Q}_{V}^{(i)},(x^{k}_{1}-x^{\star},x^{k}_{2}-x^{\star},z^{k}-z^{\star}))$
holds for all $k\in\mathbb{N}$. See Appendix˜A for the explicit expression for $\bar{Q}_{V}^{(i)}$, lifted to $\mathbb{R}^{5\times 5}$. Viewing $\bar{Q}^{(i)}_{k}$ as a quadratic form in this way will simplify the proof of Proposition˜4.3.
@@
For the sequences of Definition˜4.1 and for all $k\in\mathbb{N}$, define
@@
The quantity $I_{k}$ in Definition˜4.3 is nonnegative for all $k\in\mathbb{N}$.
@@
Let $k\in\mathbb{N}$. By Remark˜4.1 we get that each of the four inner products in $I_{k}$ is a monotonicity term; the first and third arise from the monotonicity of $\partial f$ evaluated at $(x_{1}^{k},x^{\star})$ and $(x_{1}^{k+1},x^{\star})$, while the second and fourth arise from the monotonicity of $B$ evaluated at $(x_{2}^{k},x^{\star})$ and $(x_{2}^{k+1},x^{\star})$. Therefore, $I_{k}\geq 0$. ∎
The following Lyapunov equality is the key component for establishing convergence of Algorithm˜1.
@@
With $V_{k}$ and $R_{k}$ as in Definition˜4.2 and $I_{k}$ as in Definition˜4.3, if $(\alpha,\beta,\theta)\in S^{(1)}\cup S^{(2)}$, then $V_{k}\geq 0$, $R_{k}\geq 0$, and the equality
$V_{k+1}=V_{k}-R_{k}-\theta\alpha I_{k},$ (LE)
holds, for all $k\in\mathbb{N}$.
@@
The fact that $V_{k}$ and $R_{k}$, as defined in Definition˜4.2, both are nonnegative if $(\alpha,\beta,\theta)\in S^{(1)}$ follows immediately from their respective definitions, since all coefficients in front of the norms and the Bregman-type distances are nonnegative when $(\alpha,\beta,\theta)\in S^{(1)}$ and the Bregman-type distances are nonnegative by Lemma˜2.3. For $(\alpha,\beta,\theta)\in S^{(2)}$, nonnegativity follows from the same argument applied to the alternative expressions for $V_{k}$ and $R_{k}$ of Lemma˜4.1. The Lyapunov equality (LE) is verified in Section˜A.2 and symbolically in1. ∎
@@
With the sequences of Definition˜4.1, $V_{k}$ and $R_{k}$ as in Definition˜4.2 and $I_{k}$ as in Definition˜4.3, if $(\alpha,\beta,\theta)\in S^{(1)}\cup S^{(2)}$, then
@@
    $\left(R_{k}\right)_{k\in\mathbb{N}}$ and $\left(I_{k}\right)_{k\in\mathbb{N}}$ are summable,
@@
    $\left(Q_{k}^{(1)}\right)_{k\in\mathbb{N}}$ and $\left(Q_{k}^{(2)}\right)_{k\in\mathbb{N}}$ converge.